# The Inflationary Butterfly Effect: Non-Perturbative Dynamics From Small-Scale Features

### v1

@@ Abstract
For the first time, we investigate the non-perturbative dynamics of single field inflation with a departure from slow-roll. Using simulations, we find that oscillatory features in the potential can drastically alter the course of inflation, with major phenomenological implications. In certain cases, the entire Universe gets trapped in a forever inflating de Sitter state. In others, only some regions get stuck in a false vacuum, offering an alternative channel for primordial black hole formation. Analogous to the flap of a butterfly, these results show that small-scale phenomena can have profound consequences on the evolution of the entire Universe. This demonstrates the necessity of a non-perturbative approach in the exploration of the small-scale physics of inflation, particularly in the regime relevant for gravitational-wave astronomy.
@@ I Introduction
Small-scale phenomena can profoundly influence vastly larger natural systems, as famously illustrated by the butterfly effect on the Earth’s atmosphere Lorenz (1972). This concept emerged from the first computer simulations of convective processes, which laid the foundation for chaos theory Lorenz (1963). The interplay between small and large scales can also be important in primordial cosmology. An example is Linde’s chaotic model of inflation Linde (1983, 1986), in which a tiny Planck-scale region of space-time ends up becoming the entire observable Universe. In this paper, we use simulations to explore small-scale phenomena occurring during the inflationary epoch. We demonstrate that, similarly to the flap of a butterfly, these phenomena can induce critical nonlinear consequences on the history and properties of the Universe. Inflation stands as the most natural mechanism for explaining the origin of structures in the Universe Guth (1981); Sato (1981); Linde (1987); Albrecht and Steinhardt (1982); Starobinsky (1980, 1979); Mukhanov and Chibisov (1981); Hawking (1982); Guth and Pi (1982); Starobinsky (1982); Abbott and Wise (1984). Large-scale observations, such as those of the cosmic microwave background (CMB), tightly constrain the potential of the “inflaton”, the scalar field driving inflation collaboration (2020); Akrami et al. (2020); Ade et al. (2021). By contrast, density fluctuations on smaller scales, forged in the “dark era” of inflation, remain unknown. Theoretically, four decades of work have confirmed that the sensitivity of inflation to Planck-scale physics makes it challenging to realize a long period of slow-roll inflation Baumann and McAllister (2015). Hence, inflationary scenarios deviating on small scales from the known large-scale slow-roll dynamics deserve serious scrutiny. Fortunately, the advent of gravitational-wave (GW) astronomy now provides us with a formidable opportunity to grasp that unknown cosmological epoch (see Achúcarro et al. (2022); Auclair et al. (2023); Bagui et al. (2023) for recent reviews).
@@
Deviation from slow-roll generically leads to enhanced density fluctuations. These can seed primordial black holes (PBHs) Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974); Carr (1975); Chapline (1975), acting as dark matter, and leave unique imprints in the gravitational wave background (GWB) (see the above reviews Achúcarro et al. (2022); Auclair et al. (2023); Bagui et al. (2023) and references therein). However, these large fluctuations challenge our perturbative understanding of inflation. Actually, many models generate a GWB loud enough to be detectable by future observatories, precisely when perturbative control becomes under threat Fumagalli et al. (2021); Inomata (2021); Inomata et al. (2022); Fumagalli et al. (2022); Inomata et al. (2023); Unal et al. (2023). Hence, it is crucial that theorists develop non-perturbative methods, to understand the small-scale physics of inflation and to make firm predictions to test against observations. In this work, we present the first non-perturbative study of a single-field model of inflation with enhanced density fluctuations. We consider one of the simplest and most theoretically motivated class of models, single-field inflation with oscillatory features in its potential, that we study using lattice field theory simulations. This approach has been extensively used to understand preheating after inflation (see, e.g., Khlebnikov and Tkachev (1996); Prokopec and Roos (1997); Felder and Tkachev (2008); Garcia-Bellido and Figueroa (2007); Frolov (2008); Huang (2011); Sainio (2012); Child et al. (2013); Easther et al. (2010); Lozanov and Amin (2020); Figueroa et al. (2021)), and it has more recently been extended to inflation Caravano et al. (2021, 2022, 2023); Caravano (2022); Krajewski and Turzyński (2022); Figueroa et al. (2023) (see also, e.g. , East et al. (2016); Braden et al. (2017); Clough et al. (2017, 2018) for general relativistic treatments). We demonstrate that tiny small-scale quantum fluctuations, amplified by the oscillatory feature, can drastically affect the fate of the entire Universe, sometimes preventing inflation from ending. This “inflationary butterfly effect” reveals the need for a non-perturbative treatment in the regime relevant for observations. This marks the beginning of a new era in the exploration of the Universe’s earliest epoch, reminiscent in some ways of the pioneering simulations of chaotic phenomena in Lorenz (1963).
@@ II Breakdown of perturbation theory
Inflationary potentials with oscillatory features are a well-known class of models, notably inspired by axion monodromy in string theory McAllister et al. (2010). We consider a localized oscillation:
@@
where $\phi$ is the inflaton, $V_{\rm sr}$ is a slow-roll potential and $\cal{W}$ is a smooth top-hat function that localizes the oscillatory feature between $\phi_{0}$ and $\phi_{0}+\Delta\phi$, $\mathcal{W(\phi)}=\frac{1}{4}\left(1+\tanh\left(\frac{\phi-\phi_{0}}{f}\right)\right)\left(1+\tanh\left(\frac{\phi_{0}-\phi+\Delta\phi}{f}\right)\right)$. We set $V_{\rm sr}=\bar{V}\left(1-\frac{1-n_{s}}{2}\frac{\phi^{2}}{2M_{\rm Pl}^{2}}\right)$, with $\bar{V}$ and $n_{s}$ chosen to fit the CMB-scale curvature power spectrum away from the feature. The amplitude of the oscillation is parametrized by the dimensionless quantity $\tilde{b}=2\Lambda^{4}/\dot{\phi}_{0}^{2}$ Creminelli et al. (2024). Here $\dot{\phi}_{0}^{2}=2\epsilon_{0}H_{0}^{2}M_{\rm Pl}^{2}$, with $\epsilon_{0}=\frac{\phi_{0}^{2}}{2}\left(\frac{1-n_{s}}{2}\right)^{2}$ the slow-roll parameter when the inflaton reaches the feature, and $\phi_{0}$ is a free parameter, determining which scales are affected by the feature. Throughout this study, we set $\epsilon_{0}=10^{-6}$ as a fiducial value. The oscillation frequency is $\omega=\dot{\phi}_{0}/f\equiv\alpha H_{0}$, where $H_{0}$ is the Hubble parameter at $\phi_{0}$, and $\alpha$ quantifies the number of oscillations per Hubble time. In the regime of interest with $\alpha\gg 1$, the oscillations resonate with the quantum oscillations of fluctuations inside the Hubble radius. This mechanism has been extensively studied in the regime of small-amplitude oscillation with $\tilde{b}\ll 1$, leading to so-called resonant features in the statistics of primordial fluctuations Chen et al. (2008); Flauger et al. (2010); Flauger and Pajer (2011); Behbahani et al. (2012); Duaso Pueyo and Pajer (2023); Creminelli et al. (2024). In this paper, we consider situations with $\tilde{b}\sim 1$, in which the potential is not monotonous. This regime of parameter space has only been studied recently in Inomata et al. (2023), for the three setups in Table 1.
@@ III Lattice simulations
In the simulations, we solve the fully nonlinear dynamics of the inflaton field, in a homogeneous Friedmann-Lemaître-Robertson-Walker (FLRW) metric whose dynamics is governed by the averaged density and pressure in our box. Working in such a decoupling limit by neglecting metric fluctuations is perfectly legitimate in our situations with $\epsilon\equiv-\dot{H}/H^{2}\ll 1$ Cheung et al. (2008); Behbahani et al. (2012); Creminelli et al. (2024). We also stress that, like for preheating Armendariz-Picon (2019), evolving the system using classical equations of motion is adequate in our regime of large occupation number: the large amount of particle production causing the breakdown of perturbation theory and necessitating the simulations also guarantees their validity to capture physical effects driven by classical nonlinearities.
@@ V Trapped regions as black holes
There, only some patches are stuck in the oscillatory region of the potential. From the point of view of an observer outside these trapped regions, these look as PBHs after the end of inflation. This is similar to what happens in the case of bubble nucleation due to quantum tunneling during inflation Deng and Vilenkin (2017); Garriga et al. (2016). The situation after inflation can be well described in terms of a de Sitter region embedded in a radiation- or matter-dominated (RD or MD) FLRW spacetime with a lower energy-density. While initially connected through a wormhole, the wormhole later closes, and the de Sitter regions then appear as Schwarzschild black holes for an outside observer Deng and Vilenkin (2017); Garriga et al. (2016); Inomata et al. (2022). Note that the trapping phenomenon is different from the one discussed in Deng and Vilenkin (2017); Garriga et al. (2016): it does not rely on the rare probability to tunnel to other local minima of the potential in the landscape, but instead, the minimum is lying on the classical trajectory. Such formation of PBHs from trapped regions in single-field models has been pointed out in Atal et al. (2019, 2020); Inomata et al. (2022), and is studied here for the first time with lattice simulations.
@@
Let us call ${\cal F}=V_{\rm trapped}/V_{\rm total}$ the fraction of the volume in the simulation box in which the inflaton is trapped at the end of the simulation. Due to causality, ${\cal F}$ is preserved during inflation once the trapped regions become larger than the horizon scale $1/(aH)$, which is confirmed by the simulation. The mass of PBHs produced from the trapped regions is of the same order as the horizon mass at the horizon entry of the bubbles  Deng and Vilenkin (2017); Garriga et al. (2016). Given this, we can roughly regard ${\cal F}$ as the mass fraction of PBHs at formation $\beta$ Inomata et al. (2022). The relation between $\beta$ and today’s fraction of PBHs over dark matter $f_{\rm PBH}$ depends on the PBH mass and whether they are produced during a MD or RD era. For example, for solar-mass PBHs produced during RD, $f_{\rm PBH}\sim\mathcal{O}(1)$ requires $\beta\sim 10^{-9}$ Sasaki et al. (2018). [^1] If we consider supermassive PBHs with, e.g., $M_{\text{PBH}}\sim 10^{10}\,M_{\odot}$, $f_{\text{PBH}}\simeq\mathcal{O}(1)$ corresponds to $\beta\sim 10^{-4}$, due to the proportionality $f_{\text{PBH}}\propto M_{\text{PBH}}^{-1/2}$ Sasaki et al. (2018). However, for $\mathcal{O}(10^{3})<M_{\text{PBH}}/M_{\odot}<\mathcal{O}(10^{11})$, the passage of nearby PBHs can destroy various astronomical objects, resulting in the bound $f_{\text{PBH}}\lesssim\mathcal{O}(10^{-3})$. Hence, even for supermassive PBHs, one should require $\beta\lesssim 10^{-7}$. If we consider PBHs produced during an early MD era, $\beta$ can be increased up to $\sim 10^{-6}$ for $f_{\rm PBH}\sim\mathcal{O}(1)$, with arbitrary values of the mass (see the Supplemental Material for details).
@@
Given our discussion, and considering PBHs heavier than $\mathcal{O}(10^{15})\,\text{g}$ that have not evaporated away by now  Sasaki et al. (2018), we require ${\cal F}<10^{-6}$ to avoid PBH overproduction. In case B we have ${\cal F}\sim 20\%$, which is therefore ruled out. In Fig. 3, we show the dependencies of the ratio ${\cal F}\simeq\beta$ on the values of the parameters $\tilde{b}$ and $\alpha$. We find that the value of ${\cal F}$ is extremely sensitive to variations of these parameters. Using these results, we can identify the relevant parameter range to avoid PBH overproduction, corresponding to $\tilde{b}<1.27$ or $\alpha<9$. In this regime, the simulation still exhibits an exponential tail in the PDF of the inflaton in the slow-roll part of the Universe, which we plan to investigate in future works.
@@ VI Conclusion
We focused on the case of an oscillatory feature and demonstrated that the perturbative description completely breaks down while linear perturbation theory predicts $\mathcal{P}_{\zeta}\sim 10^{-2}$. In this case, the inflaton can be trapped in local minima. The trapping can be global, affecting the whole Universe, or local, providing an alternative channel to the formation of PBHs compared to the standard collapse of high-density regions. For the latter channel, predicting the PBH abundance from the statistical properties of $\zeta$ is known to be challenging. By contrast, we were able to identify the abundance of PBHs from trapped patches directly from the simulations.
@@ I Fraction of primordial black holes produced during an early matter era
where $\Omega_{\text{DM}}$ and $\Omega_{\text{b}}$ are the energy density parameters for dark matter and baryon, and the subscript *** and eq denote, respectively, the values at the end of the eMD era, and at the time of the matter-radiation equality ($T_{\text{eq}}\simeq 8.0\times 10^{-7}\,\text{MeV}$ Inomata and Nakama (2019)). Note that for simplicity, we neglect the change in the number of degrees of freedom, which does not affect the order of magnitude of the resulting upper bound on $\beta$.

### v2

# Inflationary Butterfly Effect: Non-Perturbative Dynamics From Small-Scale Features

@@ Abstract
For the first time, we investigate the non-perturbative dynamics of single field inflation with a departure from slow-roll. Using simulations, we find that oscillatory features in the potential can drastically alter the course of inflation, with major phenomenological implications. In certain cases, the entire Universe gets trapped in a forever inflating de Sitter state. In others, only some regions get stuck in a false vacuum, offering an alternative channel for primordial black hole formation. Analogous to the flap of a butterfly, these results show that small-scale phenomena can have profound consequences on the evolution of the entire Universe. More generally, our work shows the power of simulations in the exploration of the small-scale physics of inflation, particularly in the regime relevant for gravitational-wave astronomy.
@@ I Introduction
Small-scale phenomena can profoundly influence vastly larger natural systems, as famously illustrated by the butterfly effect on the Earth’s atmosphere Lorenz (1972). This concept emerged from the first computer simulations of convective processes, which laid the foundation for chaos theory Lorenz (1963). The interplay between small and large scales can also be important in primordial cosmology. An example is Linde’s chaotic model of inflation Linde (1983, 1986), in which a tiny Planck-scale region of space-time ends up becoming the entire observable Universe. In this paper, we use simulations to explore small-scale phenomena occurring during the inflationary epoch. We demonstrate that, similarly to the flap of a butterfly, these phenomena can induce critical nonlinear consequences on the history and properties of the Universe.
Inflation stands as the most natural mechanism for explaining the origin of structures in the Universe Guth (1981); Sato (1981); Linde (1987); Albrecht and Steinhardt (1982); Starobinsky (1980, 1979); Mukhanov and Chibisov (1981); Hawking (1982); Guth and Pi (1982); Starobinsky (1982); Abbott and Wise (1984). Large-scale observations, such as those of the cosmic microwave background (CMB), tightly constrain the potential of the “inflaton”, the scalar field driving inflation collaboration (2020); Akrami et al. (2020); Ade et al. (2021). By contrast, density fluctuations on smaller scales, forged in the “dark era” of inflation, remain unknown. Theoretically, four decades of work have confirmed that the sensitivity of inflation to Planck-scale physics makes it challenging to realize a long period of slow-roll inflation Baumann and McAllister (2015). Hence, inflationary scenarios deviating on small scales from the known large-scale slow-roll dynamics deserve serious scrutiny. Fortunately, the advent of gravitational-wave (GW) astronomy now provides us with a formidable opportunity to grasp that unknown cosmological epoch (see Achúcarro et al. (2022); Auclair et al. (2023); Bagui et al. (2023) for recent reviews).
@@
Deviation from slow-roll generically leads to enhanced density fluctuations. These can seed primordial black holes (PBHs) Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974); Carr (1975); Chapline (1975), acting as dark matter, and leave unique imprints in the gravitational wave background (GWB) (see the above reviews Achúcarro et al. (2022); Auclair et al. (2023); Bagui et al. (2023) and references therein). However, these large fluctuations challenge our perturbative understanding of inflation. Actually, many models generate a GWB loud enough to be detectable by future observatories, precisely when perturbative control becomes under threat, see e.g. Fumagalli et al. (2021); Inomata (2021); Inomata et al. (2022); Fumagalli et al. (2022); Inomata et al. (2023); Unal et al. (2023); Iacconi and Mulryne (2023). Hence, it is crucial that theorists develop non-perturbative methods, to understand the small-scale physics of inflation and to make firm predictions to test against observations.
In this work, we present the first non-perturbative study of a single-field model of inflation with enhanced density fluctuations. We consider one of the simplest and most theoretically motivated class of models, single-field inflation with oscillatory features in its potential, that we study using lattice field theory simulations. This approach has been extensively used to understand preheating after inflation (see, e.g., Khlebnikov and Tkachev (1996); Prokopec and Roos (1997); Felder and Tkachev (2008); Garcia-Bellido and Figueroa (2007); Frolov (2008); Huang (2011); Sainio (2012); Child et al. (2013); Easther et al. (2010); Lozanov and Amin (2020); Figueroa et al. (2021)), and it has more recently been extended to inflation Caravano et al. (2021, 2022, 2023); Caravano (2022); Krajewski and Turzyński (2022); Figueroa et al. (2023) (see also, e.g. , East et al. (2016); Braden et al. (2017); Clough et al. (2017, 2018) for general relativistic treatments).
We demonstrate that tiny small-scale quantum fluctuations, amplified by the oscillatory feature, can drastically affect the fate of the entire Universe, sometimes preventing inflation from ending. This “inflationary butterfly effect” reveals the need for a non-perturbative treatment in the regime relevant for observations. This marks the beginning of a new era in the exploration of the Universe’s earliest epoch, reminiscent in some ways of the pioneering simulations of chaotic phenomena in Lorenz (1963).
@@ II Breakdown of perturbation theory
Inflationary potentials with oscillatory features are a well-known class of models, originally introduced in the context of primordial non-Gaussianity Chen et al. (2008), and which can be realized with axion monodromy in string theory Silverstein and Westphal (2008); McAllister et al. (2010); Flauger et al. (2010). We consider a localized oscillation:
@@
where $\phi$ is the inflaton, $V_{\rm sr}$ is a slow-roll potential and $\cal{W}$ is a smooth top-hat function that localizes the oscillatory feature between $\phi_{0}$ and $\phi_{0}+\Delta\phi$, $\mathcal{W(\phi)}=\frac{1}{4}\left(1+\tanh\left(\frac{\phi-\phi_{0}}{f}\right)\right)\left(1+\tanh\left(\frac{\phi_{0}-\phi+\Delta\phi}{f}\right)\right)$. We set $V_{\rm sr}=\bar{V}\left(1-\frac{1-n_{s}}{2}\frac{\phi^{2}}{2M_{\rm Pl}^{2}}\right)$, with $\bar{V}$ and $n_{s}$ chosen to fit the CMB-scale curvature power spectrum away from the feature.
The amplitude of the oscillation is parametrized by the dimensionless quantity $\tilde{b}=2\Lambda^{4}/\dot{\phi}_{0}^{2}$ Creminelli et al. (2024). Here $\dot{\phi}_{0}^{2}=2\epsilon_{0}H_{0}^{2}M_{\rm Pl}^{2}$, with $\epsilon_{0}=\frac{\phi_{0}^{2}}{2}\left(\frac{1-n_{s}}{2}\right)^{2}$ the slow-roll parameter when the inflaton reaches the feature, and $\phi_{0}$ is a free parameter, determining which scales are affected by the feature. Throughout this study, we set $\epsilon_{0}=10^{-6}$ as a fiducial value. The oscillation frequency is $\omega=\dot{\phi}_{0}/f\equiv\alpha H_{0}$, where $H_{0}$ is the Hubble parameter at $\phi_{0}$, and $\alpha$ quantifies the number of oscillations per Hubble time.
In the regime of interest with $\alpha\gg 1$, the oscillations resonate with the quantum oscillations of fluctuations inside the Hubble radius. This mechanism has been extensively studied in the regime of small-amplitude oscillation with $\tilde{b}\ll 1$, leading to so-called resonant features in the statistics of primordial fluctuations Chen et al. (2008); Flauger et al. (2010); Flauger and Pajer (2011); Behbahani et al. (2012); Duaso Pueyo and Pajer (2023); Creminelli et al. (2024). In this paper, we consider situations with $\tilde{b}\sim 1$, in which the potential is not monotonous. This regime of parameter space has only been studied recently in Inomata et al. (2023), for the three setups in Table 1.
@@ III Lattice simulations
In the simulations, we solve the fully nonlinear dynamics of the inflaton field, in a homogeneous Friedmann-Lemaître-Robertson-Walker (FLRW) metric whose dynamics is governed by the averaged density and pressure in our box. Working in such a decoupling limit by neglecting metric fluctuations is perfectly legitimate in our situations with $\epsilon\equiv-\dot{H}/H^{2}\ll 1$ Cheung et al. (2008); Behbahani et al. (2012); Creminelli et al. (2024). We also stress that, like for preheating Armendariz-Picon (2019), evolving the system using classical equations of motion is adequate in our regime of large occupation number: the large amount of particle production, which causes the breakdown of perturbation theory, also guarantees their validity to capture physical effects driven by classical nonlinearities.
@@ V Trapped regions as black holes
There, only some patches are stuck in the oscillatory region of the potential. From the point of view of an observer outside these trapped regions, these look as PBHs after the end of inflation. This is similar to what happens in the case of bubble nucleation due to quantum tunneling during inflation Deng and Vilenkin (2017); Garriga et al. (2016). The situation after inflation can be well described in terms of a de Sitter region embedded in a radiation- or matter-dominated (RD or MD) FLRW spacetime with a lower energy-density. While initially connected through a wormhole, the wormhole later closes, and the de Sitter regions then appear as Schwarzschild black holes for an outside observer Deng and Vilenkin (2017); Garriga et al. (2016); Inomata et al. (2022). Note that the trapping phenomenon is different from the one discussed in Deng and Vilenkin (2017); Garriga et al. (2016): it does not rely on the rare probability to tunnel to other local minima of the potential in the landscape, but instead, the minimum is lying on the classical trajectory. Such formation of PBHs from trapped regions in single-field models has been already studied in Atal et al. (2019, 2020); Inomata et al. (2022); Escrivà et al. (2023); Huang et al. (2023), both analytically and numerically, and is studied here for the first time with 3D lattice simulations.
@@
Let us call ${\cal F}=V_{\rm trapped}/V_{\rm total}$ the fraction of the volume in the simulation box in which the inflaton is trapped at the end of the simulation. Because of causality, ${\cal F}$ is preserved during inflation once the trapped regions become larger than the horizon scale $1/(aH)$, which is confirmed by the simulation. The mass of PBHs produced from the trapped regions is of the same order as the horizon mass at the horizon entry of the bubbles  Deng and Vilenkin (2017); Garriga et al. (2016). Given this, we can roughly regard ${\cal F}$ as the mass fraction of PBHs at formation $\beta$ Inomata et al. (2022). The relation between $\beta$ and today’s fraction of PBHs over dark matter $f_{\rm PBH}$ depends on the PBH mass and whether they are produced during a MD or RD era. For example, for solar-mass PBHs produced during RD, $f_{\rm PBH}\sim\mathcal{O}(1)$ requires $\beta\sim 10^{-9}$ Sasaki et al. (2018). [^1] If we consider supermassive PBHs with, e.g., $M_{\text{PBH}}\sim 10^{10}\,M_{\odot}$, $f_{\text{PBH}}\simeq\mathcal{O}(1)$ corresponds to $\beta\sim 10^{-4}$, due to the proportionality $f_{\text{PBH}}\propto M_{\text{PBH}}^{-1/2}$ Sasaki et al. (2018). However, for $\mathcal{O}(10^{3})<M_{\text{PBH}}/M_{\odot}<\mathcal{O}(10^{11})$, the passage of nearby PBHs can destroy various astronomical objects, resulting in the bound $f_{\text{PBH}}\lesssim\mathcal{O}(10^{-3})$. Hence, even for supermassive PBHs, one should require $\beta\lesssim 10^{-7}$. If we consider PBHs produced during an early MD era, $\beta$ can be increased up to $\sim 10^{-6}$ for $f_{\rm PBH}\sim\mathcal{O}(1)$, with arbitrary values of the mass (see the Supplemental Material for details).
@@
Given our discussion, and considering PBHs heavier than $\mathcal{O}(10^{15})\,\text{g}$ that have not evaporated away by now  Sasaki et al. (2018), we require ${\cal F}<10^{-6}$ to avoid PBH overproduction. In the third case considered in section IV, we have ${\cal F}\sim 20\%$, which is therefore ruled out. In Fig. 3, we show the dependencies of the ratio ${\cal F}\simeq\beta$ on the values of the parameters $\tilde{b}$ and $\alpha$. We find that the value of ${\cal F}$ is extremely sensitive to variations of these parameters. Using these results, we can identify the relevant parameter range to avoid PBH overproduction, corresponding to $\tilde{b}<1.27$ or $\alpha<9$. In this regime, the simulation still exhibits an exponential tail in the PDF of the inflaton in the slow-roll part of the Universe, which we plan to investigate in future works.
@@ VI Conclusion
We focused on the case of an oscillatory feature and demonstrated that the perturbative description completely breaks down while linear perturbation theory predicts $\mathcal{P}_{\zeta}\sim 10^{-2}$. In this case, the inflaton can be trapped in local minima. The trapping can be global, affecting the whole Universe, or local, providing an alternative channel to the formation of PBHs compared to the standard collapse of high-density regions. For the latter channel, predicting the PBH abundance from the statistical properties of $\zeta$ is known to be challenging. By contrast, we were able to estimate the abundance of PBHs from trapped patches directly from the simulations.
@@ I Fraction of primordial black holes produced during an early matter era
where $\Omega_{\text{DM}}$ and $\Omega_{\text{b}}$ are the energy density parameters for dark matter and baryon, and the subscript * and eq denote, respectively, the values at the end of the eMD era, and at the time of the matter-radiation equality ($T_{\text{eq}}\simeq 8.0\times 10^{-7}\,\text{MeV}$ Inomata and Nakama (2019)). Note that for simplicity, we neglect the change in the number of degrees of freedom, which does not affect the order of magnitude of the resulting upper bound on $\beta$.